OpenStudy (anonymous):

given f(x)=x^3+x+4 , find f^-1(1) accurate to 2 decimal places the answer is f^-1(1)=-1.21 can anyone possibly guide me threw it?

OpenStudy (dumbcow):

set f(x) = 1, then solve for x --> x^3 +x +3 = 0 try to factor out +-1 or +-3 using synthetic division

OpenStudy (raden):

it's cant be factored

OpenStudy (dumbcow):

hmm there are no rational zeroes then ...use approximation methods possible newtons method if you know calculus what class is this for?

OpenStudy (anonymous):

this is a preclac class

OpenStudy (dumbcow):

use graphing calculator :)

OpenStudy (anonymous):

im trying that now

OpenStudy (raden):

wolfram :)

OpenStudy (dumbcow):

sorry i don't know of any ways to do it by hand at a pre-calc level

OpenStudy (anonymous):

no its ok . it proabaly does want me to use a calc

OpenStudy (anonymous):

dam still confused

OpenStudy (raden):

are u sure the equation above f(x)=x^3+.... not f(x)=x^2+...

OpenStudy (anonymous):

yep

OpenStudy (raden):

hmmm.. i give up :):)

OpenStudy (anonymous):

how would i put the negative inverse in the graphing calc

OpenStudy (raden):

lok this : (im scared after looking the answer :)) http://www.wolframalpha.com/input/?i=inverse+of+y%3Dx^3%2Bx%2B4

OpenStudy (anonymous):

oh my

OpenStudy (anonymous):

lmao

OpenStudy (raden):

hahaha...

OpenStudy (anonymous):

thanks man i appreciate you taking away your time to help me

OpenStudy (raden):

you're welcome